40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8362     Entry time: Thu Mar 28 00:31:11 2013
Author: Jenne 
Type: Update 
Category: SUS 
Subject: PRMI optics' oplev servos tuned 

[Rana, Gabriele, Jenne, Jamie, Lisa, Rana]

We have tuned the oplev servos for PRM, BS, ITMX, ITMY.  For each, we measured the servo transfer function.  Most had a UGF ~ 3Hz.  For those, we increased the gain by a factor of 2, and engaged the 3.3Hz resonant gains. The other case, such as PRM yaw, the gain was already okay, we just needed to engage the resonant gain.  We also checked the new phase margin, and for some of them switched the elliptic lowpass to 50Hz rather than 30 or 35.

Before and afters:

PRM_OLPIT_tuning_27Mar2013_RedIsAfter.pdf

PRM_OLYAW_tuning_27Mar2013_RedIsAfter.pdf

bs_pit.pdf

bs_yaw.pdf

ITMY_OLPIT_tuning_27Mar2013_RedIsAfter.pdf

ITMY_OLYAW_tuning_27Mar2013_RedIsAfter.pdf

itmx_pit1.pdf

itmx_yaw.pdf

We need to, as a last check, look at the spectra before and after to ensure that no modes (like bounce or roll) are newly excited.

ELOG V3.1.3-